Kris808:

Which set of line segments could create a right triangle? 24, 30, 35 12, 18, 30 18, 24, 30 18, 24, 35

ramen:

Try using the pathagrorem thereom

ramen:

a^2+b^2=c^2

Kris808:

O-O you really are a human cauculator.

ramen:

A and B are the 2 smallest sides and c is the biggest side.

ramen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Kris808 O-O you really are a human cauculator. \(\color{#0cbb34}{\text{End of Quote}}\) lol thanks

ramen:

Do you know what to do now?

Kris808:

Yes

Kris808:

Thank you

umm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen a^2+b^2=c^2 \(\color{#0cbb34}{\text{End of Quote}}\) \[a ^{2}+b ^{2}=c ^{2}\] Easier understanding.

Kris808:

o-o thanks to you too i guess

umm:

Anytime. (: